What is the mass of a 3.90-?Ci 146C source? The half-life of 146C is 5730 yr.

Express your answer to three significant figures and include the appropriate units.

please help. I have no idea what im doing on this.
Answer #1


activity, A=3.9*10^-6 Ci


A=3.9*10^-6*(3.7*10^10)


A=1.443*10^5 decays/sec


use,


no of atoms,


N=A*t1/2/(ln(2))


N=1.443*10^5*5730*365*24*60*60/(ln(2))


=3.762*10^16

mass of the atom, m=N/NA


=3.762*10^16/(6.02*10^23)


=6.25*10^-8 kg


total mass, M=14*6.25*10^-8


M=8.75*10^-7 kg


or

M=8.78*10^-4 g